laineyluvsu:

What is the measure of 1 if 1 is (4x - 4) and 2 is (4x) ?

InPinkLuv:

alright so we got a 1= (4x - 4) and 2 = 4x. but you want to try to find the value of 1 in terms of x, right?

laineyluvsu:

@inpinkluv wrote:
alright so we got a 1= (4x - 4) and 2 = 4x. but you want to try to find the value of 1 in terms of x, right?
Yes

InPinkLuv:

since we got 2=4x, we can substitute that equation into a 1.

InPinkLuv:

so let just say 2 = 4x, we can rewrite it as 1

InPinkLuv:

so it would be 1= (2 - 4), so 1= -2.

InPinkLuv:

and that's the measure of the 1, fam so yeah :3

InPinkLuv:

hopefully this helps :)

fwval:

the measure of 1 is -2, given that 1 is (4x - 4) and 2 is (4x).

fwval:

@inpinkluv wrote:
so it would be 1= (2 - 4), so 1= -2.
or this